Need assistance solving problem.
. > Moving to another question will save this response. Question1 What is the proportion of scores in a normal distribution between the mean(z-0.00) and z +0.527 0 0.52 Report your answer at four decimal places. will save this response. a) Moving to another q
Answer #1

To find the proportion between the mean that is z = 0 and z = 0.52, use z distribution table

The table gives the left side area to the z

Proportion between 0 to 0.52 means P(0 < z < 0.52)

which can be written as

P(0 < z < 0.52) = P(z < 0.52) - P(z < 0)

The probability of z score for 0 using z table is 0.5000 and the probability of z score for 0.52 is 0.6985

Therefore, P(0 < z < 0.52) = 0.6985 - 0.5000 = 0.1985

Therefore, the proportion of scores in a normal distribution between the mean (z = 0) and z = 0.52 is 0.1985